Oil flows through a pipeline which contracts from 450mm diameter at A to 300mm diameter at B and then forks, one branch being 150mm diameter discharging at C and the other branch 225mm diameter discharging at D. if the velocity at A is 1.8 m/s and the velocity at D is 3.6 m/s, what will be the discharges at C & D and the velocity at B & C

1
Expert's answer
2022-04-11T16:55:35-0400

"Q_C=vS_C=0.03~\frac{m^3}s,"

"Q_D=vS_D=0.07~\frac{m^3}s,"

"v_B=v\frac{S_B}{S_B+S_C}=1.3~\frac ms,"

"v_C=v\frac{S_C}{S_B+S_C}=0.8~\frac ms."